# Stabilizing the Splits through Minimax Decision Trees

### v2

@@ B.2 Proofs of results in Section 3
for some (random) positive integers $\xi_{k,i},\leavevmode\nobreak\ i\in[d]$ depending on the samples from $\Sigma$. See also (61) below. The intuition in proving Theorem 9 is that we keep (57) as an upper bound, instead of (54). Then, with enough randomness injected into the splitting dimensions (and hence the random variables $\{\xi_{k,i}\}$), the right-hand side of (57) is small if averaged over all trees in the forest, as shown by the next lemma.
@@
In the following, we use $S$ to denote a random variable with law $\Sigma$. We first establish an upper bound of
@@
Here, the outer expectation is with respect to the randomness from the empirical measure $\mathbb{E}_{n}$ (the law of $(\mathbf{X},Y)$ is fixed in the settings of Theorems 2 and 6), and the inner expectation is with respect to the joint law of $(\mathbf{X},Y,\{M_{k}^{S}\})$. In other words, we trivially bound the risk of the average by the average of the risks among the $n$ trees.
@@
    $\xi_{k,\zeta_{k}}=\max\{\xi^{\prime}_{k-1,\zeta_{k}},\xi^{\prime\prime}_{k-1,\zeta_{k}}\};\quad\xi_{k,i}=\xi^{\prime}_{k-1,i}+\xi^{\prime\prime}_{k-1,i},\leavevmode\nobreak\ i\neq\zeta_{k}.$
@@
    $\sum_{j\in J_{k}}\Delta g_{j,i}^{\ell}$ $\leqslant\sum_{\begin{subarray}{c}j\in J_{k}\\ j\succ u\end{subarray}}\Delta g_{j,i}^{\ell}+\sum_{\begin{subarray}{c}j\in J_{k}\\ j\succ v\end{subarray}}\Delta g_{j,i}^{\ell}$
@@
  • For dimension $\zeta_{k}$, we again use the fact that the variation $\sum_{j\in J_{k}}\Delta g_{j,\zeta_{k}}^{\ell}$ is at most the sum of variations among the descendants of $u,v$, but the total variation $\left\lVert g_{\zeta_{k}}\right\rVert_{\mathrm{TV}}$ is already decomposed into two, say $\left\lVert g_{\zeta_{k}}^{(i)}\right\rVert_{\mathrm{TV}},\,i=1,2$, corresponding to $u,v$. Therefore, by inductive hypothesis,
    $\sum_{j\in J_{k}}\Delta g_{j,\zeta_{k}}^{\ell}$ $\leqslant\sum_{\begin{subarray}{c}j\in J_{k}\\ j\succ u\end{subarray}}\Delta g_{j,\zeta_{k}}^{\ell}+\sum_{\begin{subarray}{c}j\in J_{k}\\ j\succ v\end{subarray}}\Delta g_{j,\zeta_{k}}^{\ell}$
@@
is carried until (B.34) of Klusowski and Tian (2024). The term $2^{k}U^{2}/N$ is absorbed by the term $U^{4}2^{k}\log(Nd)/N$ therein. The other term $2^{k}U^{2}d/(Nn^{1/3})$ is carried into (B.35), and is bounded by $C2^{k}(\log N)d/(Nn^{1/3})$ under the choice of $U=\left\lVert g^{*}\right\rVert_{\infty}+2\sigma\sqrt{2\log N}$ at the end of that proof, where $C>0$ may depend on $\left\lVert g^{*}\right\rVert_{\infty}$ and $\sigma$. This concludes (26). ∎
@@ B.3 Proofs of results in Section 4
Suppose that $\mathbf{X}$ is marginally atomless and $j\in[d]$. Then the function $\varphi_{L}$ is non-decreasing and continuous, and $\varphi_{R}$ is non-increasing and continuous. Moreover, under the (cyclic) MinimaxSplit construction,
@@
We may assume that $\sup\operatorname{supp}X-\inf\operatorname{supp}X$, otherwise the claim is trivial. Given the marginal law of $X$, the right-hand side of (62) is minimized if $X$ and $Y$ are comonotonic. This is achieved by setting, with $a=\inf\operatorname{supp}X\leqslant\sup\operatorname{supp}X=b$, that $p=\mathbb{P}(X=b,Y=1)=1-\mathbb{P}(X=a,Y=-1)$, for some $p\in[0,1]$. It remains to show the deterministic inequality
$\kappa_{a,b}(p):=h(p)+(p\log(1+e^{-b})+(1-p)\log(1+e^{a}))+\frac{1}{4}|b-a|\geqslant 0.$
@@
We first recall the notations from the proof of Theorem 2. Denote by $\pi_{k}=\{I_{j}\}_{j\in J_{k}}$ the partition of $\mathbb{R}^{d}$ formed at level $k$ from the cyclic MinimaxSplit construction. Note that $M_{k}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}=y_{k,j}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}$, where $y_{k,j}=\log(\mathbb{P}(Y=1\mid\mathbf{X}\in I_{j})/\mathbb{P}(Y=-1\mid\mathbf{X}\in I_{j}))$. Define $g_{j}=g|_{I_{j}}$, so that $\left\lVert g\right\rVert_{\mathrm{TV}}=\sum_{j}\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$. Recall that $\Delta g_{j}=\sup g_{j}-\inf g_{j}\leqslant\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
@@
$V_{j}=\min\Big{\{}\mathbb{E}[\log(1+e^{-YM_{k}}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}],\frac{p_{j}\Delta g_{j}}{4}\Big{\}}.$
We have by construction and (64) that
@@
because adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most $\log N/N$ (which can be seen from the extreme case with a sample $1$ added to samples of 1-1). As a consequence, if we denote by
@@
$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{\log N}{N},\leavevmode\nobreak\ k\geqslant 0;\quad u_{0}=H(Y).$
@@
$\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A))=u_{k}\leqslant 2^{-k}H(Y)+\frac{2\log N}{N}.$ (70)
It is also easy to check that (70) is also satisfied for the case $u_{0}<2\log N/N$, that is, (70) holds in general. Replacing (68) by (70) leads to
$\mathbb{E}[(Y-M_{k})^{2}]\leqslant\inf_{g\in\mathcal{G}}\bigg{(}\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}\Big{(}\left\lVert g\right\rVert_{\mathrm{TV}}(1+2^{k+1}\frac{\log N}{N})\Big{)}^{1/2}\bigg{)}.$
@@
$\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}\Big{(}2^{k+1}\frac{\log N}{N}\Big{)}^{1/2}\leqslant\frac{\left\lVert g\right\rVert_{\mathrm{TV}}}{2}+\frac{2^{k}\log N}{N},$
@@ Appendix C Additional example for anti-symmetry-breaking preference
In this appendix, we provide an empirical version of Example 8, showing that the ASBP persists even when samples are involved and the symmetry is not perfect.
@@
Define $f(\mathbf{x})=x_{1}+|x_{2}|,\leavevmode\nobreak\ \mathbf{x}=(x_{1},x_{2})\in[-1,1]^{2}$, which is Lipschitz and bounded. Let $\mathbb{P}_{*}$ denote the law of $\mathbf{X}$ in Example 8 and $\mathbb{P}_{n}$ denote the empirical law of $\{(\mathbf{X}_{i},Y_{i})\}_{1\leqslant i\leqslant n}$, where each $\mathbf{X}_{i}$ is sampled from $\mathbb{P}_{*}$ and $Y_{i}=f(\mathbf{X}_{i})$. Denote by $\mathcal{R}$ the set of all axis-aligned rectangles in $1,1]^{2}$. It follows from standard arguments in van der Vaart and Wellner (2013) that the class $\{f{\mathbbm{1}}_{R}\}_{R\in\mathcal{R}}$ is a Donsker class, so with $m_{R}=\mathbb{E}^{\mathbb{P}_{*}}[f(\mathbf{X})\mid\mathbf{X}\in R]$, we have $\sup_{R\in\mathcal{R}}|\mathbb{E}^{\mathbb{P}_{n}}[f(\mathbf{X})\mid\mathbf{X}\in R]-m_{R}|=O_{\mathbb{P}}(1/\sqrt{n})$. Here, we use the notation $\xi_{n}=O_{\mathbb{P}}(\alpha_{n})$ if the sequence $\{\xi_{n}/\alpha_{n}\}$ is tight. Let $g_{R}(\mathbf{x})=(f(\mathbf{x})-m_{R})^{2}$ for $R\in\mathcal{R}$. The class $\{g_{R}{\mathbbm{1}}_{R}\}_{R\in\mathcal{R}}$ is also Donsker. For $R\in\mathcal{R}$, let $V_{R}$ denote the empirical risk on $R$, i.e.,
$V_{R}=\frac{1}{n}\sum_{i=1}^{n}((f(\mathbf{X}_{i})-\mathbb{E}^{\mathbb{P}_{n}}[f(\mathbf{X})\mid\mathbf{X}\in R])^{2}{\mathbbm{1}}_{\mathbf{X}_{i}\in R}).$
It follows from the Donsker property and triangle inequality that
@@
Also denote by $m(p)=\mathbb{E}[(M-1){\mathbbm{1}}_{\{M\geqslant 1\}}/M]$ where $M\buildrel\mathrm{law}\over{\sim}\mathrm{Bin}(n,p)$. A standard computation (by first conditioning on the number of samples in the rectangle) yields that
@@
In the following, the constants $C$ (depending on $\varepsilon$ only) may vary from line to line. Note that m(p)=1O(1/(np))m(p)=1-O(1/(np)) jointly as $np\to\infty$ (see e.g., Corollary 1 of Wang et al. (2008)). The following results hold if $b-a\geqslant C/n^{1/6}$ for $n$ large enough.
@@
Appendix D Further numerics
D.1 Performance metrics

### v3

@@ B.2 Proofs of results in Section 3
for some (random) positive integers $\xi_{k,i},\penalty 10000\ i\in[d]$ depending on the samples from $\Sigma$. See also (62) below. The intuition in proving Theorem 10 is that we keep (58) as an upper bound, instead of (54). Then, with enough randomness injected into the splitting dimensions (and hence the random variables $\{\xi_{k,i}\}$), the right-hand side of (58) is small if averaged over all trees in the forest, as shown by the next lemma.
@@
In the following, we use $S$ to denote a random variable with law $\Sigma$. We first establish an upper bound for
@@
Here, the outer expectation is with respect to the randomness from the empirical measure $\mathbb{E}_{n}$ (the law of $(\mathbf{X},Y)$ is fixed in the settings of Theorems 3 and 7), and the inner expectation is with respect to the joint law of $(\mathbf{X},Y,\{M_{k}^{S}\})$. In other words, we trivially bound the risk of the average by the average of the risks among the $n$ trees.
@@
    $\xi_{k,\zeta_{k}}=\max\{\xi^{\prime}_{k-1,\zeta_{k}},\xi^{\prime\prime}_{k-1,\zeta_{k}}\};\quad\xi_{k,i}=\xi^{\prime}_{k-1,i}+\xi^{\prime\prime}_{k-1,i},\penalty 10000\ i\neq\zeta_{k}.$
@@
    $\sum_{j\in J_{k}^{\ell}}\Delta g_{j,i}^{\ell}$ $\leqslant\sum_{\begin{subarray}{c}j\in J_{k}^{\ell}\\ j\succ u\end{subarray}}\Delta g_{j,i}^{\ell}+\sum_{\begin{subarray}{c}j\in J_{k}^{\ell}\\ j\succ v\end{subarray}}\Delta g_{j,i}^{\ell}$
@@
  • For dimension $\zeta_{k}$, we again use the fact that the variation $\sum_{j\in J_{k}^{\ell}}\Delta g_{j,\zeta_{k}}^{\ell}$ is at most the sum of variations among the descendants of $u,v$, but the total variation $\left\lVert g_{\zeta_{k}}\right\rVert_{\mathrm{TV}}$ is already decomposed into two, say $\left\lVert g_{\zeta_{k}}^{(i)}\right\rVert_{\mathrm{TV}},\,i=1,2$, corresponding to $u,v$. Therefore, by inductive hypothesis,
    $\sum_{j\in J_{k}^{\ell}}\Delta g_{j,\zeta_{k}}^{\ell}$ $\leqslant\sum_{\begin{subarray}{c}j\in J_{k}^{\ell}\\ j\succ u\end{subarray}}\Delta g_{j,\zeta_{k}}^{\ell}+\sum_{\begin{subarray}{c}j\in J_{k}^{\ell}\\ j\succ v\end{subarray}}\Delta g_{j,\zeta_{k}}^{\ell}$
@@
is carried until (B.34) of Klusowski and Tian (2024). The term $2^{k}U^{2}/N$ is absorbed by the term $U^{4}2^{k}\log(Nd)/N$ therein. The other term $2^{k}U^{2}d/(Nn^{1/3})$ is carried into (B.35), and is bounded by $C2^{k}(\log N)d/(Nn^{1/3})$ under the choice of $U=\left\lVert g_{*}\right\rVert_{\infty}+2\sigma\sqrt{2\log N}$ at the end of that proof, where $C>0$ may depend on $\left\lVert g_{*}\right\rVert_{\infty}$ and $\sigma$. This concludes (26). ∎
To prove Theorem 11, we first record the expected oscillation estimate that replaces the deterministic cyclic counting argument.
Lemma B.4.
Assume the setting of Definition 3.1, and let
$g(\mathbf{x})=\sum_{j\in S}g_{j}(x_{j})\in\mathcal{G}_{S}.$
Then for every $k\geqslant 0$,
$\mathbb{E}\left[\sum_{A\in\pi_{k}^{\Xi}}\left(\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x})\right)\right]\leqslant 2^{k}e^{-\rho k/2}\sum_{j\in S}\|g_{j}\|_{\mathrm{TV}}.$
Consequently,
$\mathbb{E}\left[\sum_{A\in\pi_{k}^{\Xi}}\left(\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x})\right)\right]\leqslant 2^{k}e^{-\rho k/2}\|g\|_{\mathrm{TV}}.$
Proof.
Recall the notation from Lemma B.2. For $j\in S$, define
$T_{j}^{\Xi}(k):=\sum_{A\in\pi_{k}^{\Xi}}\mathrm{osc}_{I_{j}(A)}(g_{j}).$
As in the proof of Lemma B.2,
$\sup_{A}g-\inf_{A}g\leqslant\sum_{j\in S}\mathrm{osc}_{I_{j}(A)}(g_{j}),$
so it suffices to bound $\mathbb{E}[T_{j}^{\Xi}(k)]$.
Fix $j\in S$. Conditional on $\pi_{k}^{\Xi}$, each splittable cell $A\in\pi_{k}^{\Xi}$ is split along coordinate $j$ with conditional probability at least $\rho$. If $A$ is split along coordinate $j$, then the sum of the oscillations of $g_{j}$ over the two children is at most $\mathrm{osc}_{I_{j}(A)}(g_{j})$. If $A$ is split along a coordinate different from $j$, then the $j$th interval is duplicated, so the sum of the oscillations over the two children is exactly $2\mathrm{osc}_{I_{j}(A)}(g_{j})$.
Therefore,
$\mathbb{E}\!\left[T_{j}^{\Xi}(k+1)\,\middle|\,\pi_{k}^{\Xi}\right]\leqslant\bigl(\rho\cdot 1+(1-\rho)\cdot 2\bigr)T_{j}^{\Xi}(k)=(2-\rho)T_{j}^{\Xi}(k).$
Taking expectations and iterating,
$\mathbb{E}[T_{j}^{\Xi}(k)]\leqslant(2-\rho)^{k}T_{j}^{\Xi}(0).$
Since $2-\rho=2(1-\rho/2)\leqslant 2e^{-\rho/2}$ and $T_{j}^{\Xi}(0)\leqslant\|g_{j}\|_{\mathrm{TV}}$, we obtain
$\mathbb{E}[T_{j}^{\Xi}(k)]\leqslant 2^{k}e^{-\rho k/2}\|g_{j}\|_{\mathrm{TV}}.$
Summing over $j\in S$ proves the first claim, and taking the infimum over additive representations proves the second. ∎
Proof of Theorem 11.
Fix a realization of the randomization $\Xi$. Conditional on $\Xi$, the tree is deterministic, and the part of the proof of Theorem 3 that leads to (45) goes through verbatim, with a suitable adaptation of the upper bound (54) for the oscillation term. More precisely, for every $g\in\mathcal{G}_{S}$, the conditional predictor $M_{k}^{\Xi}$ satisfies
$\mathbb{E}\!\left[(Y-M_{k}^{\Xi})^{2}\,\middle|\,\Xi\right]\leqslant\,$ $(1+\delta)\mathbb{E}[(Y-g(\mathbf{X}))^{2}]+(1+\delta^{-1})(\mathrm{Var}(Y)\Theta_{k}(\Xi;g))^{2/3},$ (63)
where
$\Theta_{k}(\Xi;g):=2^{-k}\sum_{A\in\pi_{k}^{\Xi}}\left(\sup_{\mathbf{x}\in A}g(\mathbf{x})-\inf_{\mathbf{x}\in A}g(\mathbf{x})\right).$
Taking expectations over $\Xi$ and using Jensen’s inequality for the concave map $x\mapsto x^{2/3}$ yields
$\mathbb{E}\!\left[\Theta_{k}(\Xi;g)^{2/3}\right]\leqslant\left(\mathbb{E}[\Theta_{k}(\Xi;g)]\right)^{2/3}.$
By Lemma B.4,
$\mathbb{E}[\Theta_{k}(\Xi;g)]\leqslant 2^{-k}\cdot 2^{k}e^{-\rho k/2}\|g\|_{\mathrm{TV}}=e^{-\rho k/2}\|g\|_{\mathrm{TV}}.$
Hence
$\mathbb{E}\!\left[\Theta_{k}(\Xi;g)^{2/3}\right]\leqslant e^{-\rho k/3}\|g\|_{\mathrm{TV}}^{2/3}.$
Substituting this estimate into (63) and then following step I of the proof of Theorem 3 gives (27). ∎
@@ B.3 Proofs of results in Section 4
Suppose that $\mathbf{X}$ is marginally atomless and $j\in[d]$. Then the function $\varphi_{L}$ is non-decreasing and continuous, and $\varphi_{R}$ is non-increasing and continuous. Moreover, denote by $\pi_{k}=\{I_{j}\}_{j\in J_{k}}$ the partition of $\mathbb{R}^{d}$ formed at level $k$ from the cyclic MinimaxSplit construction. Then
@@
We may assume that $\sup\operatorname{supp}X-\inf\operatorname{supp}X<\infty$, otherwise the claim is trivial. Given the marginal law of $X$, the right-hand side of (64) is minimized if $X$ and $Y$ are comonotonic. Moreover, with $\sup\operatorname{supp}X$ and $\inf\operatorname{supp}X$ fixed, the right-hand side of (64) is minimized when we have the equivalence of events $\{X=\sup\operatorname{supp}X\}=\{Y=1\}$ and $\{X=\inf\operatorname{supp}X\}=\{Y=-1\}$. This is achieved by setting, with $a=\inf\operatorname{supp}X\leqslant\sup\operatorname{supp}X=b$, that $p=\mathbb{P}(X=b,Y=1)=1-\mathbb{P}(X=a,Y=-1)$, for some $p\in[0,1]$. It remains to show the deterministic inequality
$\kappa_{a,b}(p):=-h(p)+(p\log(1+e^{-b})+(1-p)\log(1+e^{a}))+\frac{1}{4}|b-a|\geqslant 0.$
@@
We first recall the notations from the proof of Theorem 3. Denote by $\pi_{k}=\{I_{j}\}_{j\in J_{k}}$ the partition of $\mathbb{R}^{d}$ formed at level $k$ from the cyclic MinimaxSplit construction. Note that $M_{k}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}=y_{k,j}{\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}$, where $y_{k,j}=\log(\mathbb{P}(Y=1\mid\mathbf{X}\in I_{j})/\mathbb{P}(Y=-1\mid\mathbf{X}\in I_{j}))$. Define $g_{j}=g|_{I_{j}}$ and recall that $\Delta g_{j}=\sup g_{j}-\inf g_{j}\leqslant\left\lVert g_{j}\right\rVert_{\mathrm{TV}}$.
@@
$V_{j}=\min\Big\{\mathbb{E}[\log(1+e^{-YM_{k}}){\mathbbm{1}}_{\{\mathbf{X}\in I_{j}\}}],\frac{p_{j}\Delta g_{j}}{4}\Big\},$
where $p_{j}:=\mathbb{P}(\mathbf{X}\in I_{j})$. We have by construction and (66) that
@@
because adding an atom of weight $\leqslant 1/N$ to a node increases the risk by at most
$\frac{1}{N}\,\log N+\Big(1-\frac{1}{N}\Big)\log\Big(\frac{1}{1-\frac{1}{N}}\Big)\leqslant\frac{1+\log N}{N},$
which can be seen from the extreme case with a sample $1$ added to samples of 1-1. As a consequence, if we denote by
@@
$u_{k+1}\leqslant\frac{u_{k}}{2}+\frac{\log(eN)}{N},\penalty 10000\ k\geqslant 0;\quad u_{0}=H(Y).$
@@
$\max_{A\in\pi_{k}}\mathbb{P}(\mathbf{X}\in A)\,h(\mathbb{P}(Y=1\mid\mathbf{X}\in A))=u_{k}\leqslant 2^{-k}H(Y)+\frac{2\log(eN)}{N}.$ (72)
It is also easy to check that (72) is also satisfied for the case $u_{0}<2\log(eN)/N$, that is, (72) holds in general. Replacing (70) by (72) leads to
$\mathbb{E}[\log(1+e^{-YM_{k}})]\leqslant\inf_{g\in\mathcal{G}}\bigg(\mathbb{E}[\log(1+e^{-Yg(\mathbf{X})})]+\sqrt{\frac{\log 2}{4}}\,2^{-\lfloor k/d\rfloor/2}\Big(\left\lVert g\right\rVert_{\mathrm{TV}}(1+2^{k+1}\frac{\log(eN)}{N})\Big)^{1/2}\bigg).$
@@
$\left\lVert g\right\rVert_{\mathrm{TV}}^{1/2}\Big(2^{k+1}\frac{\log N}{N}\Big)^{1/2}\leqslant\frac{\left\lVert g\right\rVert_{\mathrm{TV}}}{2}+\frac{2^{k}\log(eN)}{N},$
@@
Proof of Theorem 14.
The proof follows in exactly the same way as Theorem 4.3 of Klusowski and Tian (2024) in the C4.5 case, as the arguments therein do not depend on how the tree is constructed, as long as the splitting rule is axis-aligned. ∎
@@ Appendix C Additional example for anti-symmetry-breaking preference
In this appendix, we provide an empirical version of Example 9, showing that the ASBP persists even when samples are involved (meaning that the symmetry is not perfect under the empirical measure).
@@
Define $f(\mathbf{x})=x_{1}+|x_{2}|,\penalty 10000\ \mathbf{x}=(x_{1},x_{2})\in[-1,1]^{2}$, which is Lipschitz and bounded. Let $\mathbb{P}_{*}$ denote the law of $\mathbf{X}$ in Example 9 and $\mathbb{P}_{n}$ denote the empirical law of $\{(\mathbf{X}_{i},Y_{i})\}_{1\leqslant i\leqslant n}$, where each $\mathbf{X}_{i}$ is sampled from $\mathbb{P}_{*}$ and $Y_{i}=f(\mathbf{X}_{i})$. Denote by $\mathcal{R}$ the set of all axis-aligned rectangles in $1,1]^{2}$.
Since $\left\lVert f\right\rVert_{\infty}\leqslant 2$ and the set of axis-aligned rectangles forms a VC class, the class
$\mathcal{H}:=\{(f-m)^{2}\mathbf{1}_{R}:\ R\in\mathcal{R},\ m\in[-2,2]\}$
is a bounded Donsker class. For
$T_{R}:=\inf_{m\in[-2,2]}\mathbb{E}^{\mathbb{P}_{*}}[(f(\mathbf{X})-m)^{2}\mathbf{1}_{\{\mathbf{X}\in R\}}],\qquad V_{R}:=\inf_{m\in[-2,2]}\mathbb{E}^{\mathbb{P}_{n}}[(f(\mathbf{X})-m)^{2}\mathbf{1}_{\{\mathbf{X}\in R\}}],$
we therefore have
$\sup_{R\in\mathcal{R}}|V_{R}-T_{R}|\leqslant\sup_{h\in\mathcal{H}}|\mathbb{E}^{\mathbb{P}_{*}}[h]-\mathbb{E}^{\mathbb{P}_{n}}[h]|=O_{\mathbb{P}}\Big(\frac{1}{\sqrt{n}}\Big).$
Here, we use the notation $\xi_{n}=O_{\mathbb{P}}(\alpha_{n})$ if the sequence $\{\xi_{n}/\alpha_{n}\}$ is tight. Moreover, for each $R\in\mathcal{R}$, if $M_{R}:=\sum_{i=1}^{n}\mathbf{1}_{\{\mathbf{X}_{i}\in R\}}$, then conditional on $M_{R}=m\geqslant 1$,
$\mathbb{E}[V_{R}\mid M_{R}=m]=\frac{m-1}{n}\mathrm{Var}(f(\mathbf{X})\mid\mathbf{X}\in R),$
while $V_{R}=0$ when $m=0$. Hence
$0\leqslant T_{R}-\mathbb{E}[V_{R}]=\frac{\mathbb{P}(M_{R}\geqslant 1)}{n}\mathrm{Var}(f(\mathbf{X})\mid\mathbf{X}\in R)\leqslant\frac{\|f\|_{\infty}^{2}}{n}\leqslant\frac{4}{n}.$
By the triangle inequality,
@@
Also denote by $m(p)=\mathbb{E}[(M-1){\mathbbm{1}}_{\{M\geqslant 1\}}/M]$ where $M\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Bin}(n,p)$. A standard computation (by first conditioning on the number of samples in the rectangle) yields that
@@
In the following, the constants $C$ (depending on $\varepsilon$ only) may vary from line to line. Note that m(p)=1O(1/(np))m(p)=1-O(1/(np)) jointly as $np\to\infty$ (see e.g., Corollary 1 of Wang and others (2008)). The following results hold if $b-a\geqslant C/n^{1/6}$ for $n$ large enough.
@@
Appendix D Additional synthetic experiments for the classification setting
In this appendix, we complement Section 4 with additional synthetic experiments that isolate the behavior of the classification splitting rules in one dimension. Throughout, we consider $Y\in\{\pm 1\}$ and write
$\eta(x)\coloneqq\mathbb{P}(Y=1\mid X=x),\qquad g_{*}(x)\coloneqq\log\Big(\frac{\eta(x)}{1-\eta(x)}\Big).$
We compare three axis-aligned tree constructions: (i) entropy-CART, which minimizes the weighted sum of child entropies; (ii) Gini-CART, which minimizes the weighted sum of child Gini impurities; [^10] The criterion $h(p)=-p\log p-(1-p)\log(1-p)$ was used for entropy-CART, while Gini-CART uses the Gini impurity criterion 2p(1p)2p(1-p). and (iii) the entropy-based MinimaxSplit rule from Section 4, which minimizes the maximum of the two weighted child entropies. Since $d=1$ in all experiments below, MinimaxSplit and cyclic MinimaxSplit coincide. Unless otherwise stated, each tree is grown with minimum leaf size $8$, training sample size $250$, and depth budget $k\in\{1,\dots,8\}$.
D.1 Piecewise signal: predictive risk as a function of depth
We first consider a piecewise-constant Bayes log-odds,
$g_{*}(x)=\begin{cases}2,&0\leqslant x<0.30,\\ -1.5,&0.30\leqslant x<0.60,\\ 1,&0.60\leqslant x\leqslant 1,\end{cases}\qquad X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1),$
so that $\eta(x)=e^{g_{*}(x)}/(1+e^{g_{*}(x)})$ is piecewise constant with two change-points. The left panel of Figure 12 reports the expected log-loss, which is the empirical analogue of the surrogate risk in (30) and (32). The middle panel reports the corresponding misclassification error.
Two features of Figure 12 are worth emphasizing. First, entropy-CART and Gini-CART achieve the best shallow-tree performance: at depth $k=2$ they attain an expected log-loss of around $0.52$, whereas MinimaxSplit is still near $0.54$. This is consistent with the fact that the CART rules are more aggressive in optimizing the immediate impurity decrease. Second, once the depth exceeds the first few levels, the ranking reverses. From depth $k=3$ onward, MinimaxSplit has the smallest expected log-loss, and the gap becomes substantial by depths $k=5,6,7,8$. In particular, the MinimaxSplit curve stabilizes near $0.541$, while the two CART baselines drift upward toward $0.589$. Thus, on this piecewise signal, the worst-child criterion trades some initial greediness for a substantially more stable partition at moderate and large depths.
The middle panel shows that the advantage is more pronounced for cross-entropy than for the 011 loss. This is also in line with Section 4: the primary theoretical control is on $\mathbb{E}[\log(1+e^{-YM_{k}})]$, and the excess classification risk is obtained only after the calibration step of Theorem 14. Empirically, the misclassification curves are close after depth $k=3$, whereas the cross-entropy curves remain clearly separated. The right panel shows that MinimaxSplit reaches a larger number of terminal cells by depth $5$, but, as we explain next, these leaves are much more evenly populated.
Figure 12: Synthetic one-dimensional classification experiment with piecewise Bayes log-odds. Left: expected log-loss versus depth. Middle: test misclassification versus depth. Right: average number of leaves versus depth. In dimension $d=1$, the entropy-based MinimaxSplit rule coincides with cyclic MinimaxSplit. The main empirical advantage appears in the surrogate risk: after a short shallow-tree disadvantage, MinimaxSplit attains a lower and markedly more stable expected log-loss than entropy-CART and Gini-CART.
D.2 Leaf-balance as the geometric mechanism
Figure 13 makes the geometric mechanism behind the previous risk comparison explicit. The left panel plots the mean number of samples per leaf on a logarithmic scale, and the right panel plots the standard deviation of the leaf sizes, again on a logarithmic scale. The mean leaf sizes are broadly comparable once the depth is moderate, but the dispersion is dramatically different. For example, at depth $k=4$ the standard deviation of the leaf sizes is about $4.9$ for MinimaxSplit, compared with roughly $13$ for the two CART baselines; at depth $k=6$ the corresponding values are about $3.3$ versus about 6.5.
This is the empirical signature of the worst-child criterion. Entropy-CART and Gini-CART reduce the average impurity, and therefore can continue to benefit from highly unbalanced splits as long as one child becomes pure enough. By contrast, MinimaxSplit penalizes any split that leaves one child both large and highly impure. The resulting trees are therefore more regular in the sense that the sample sizes of the terminal cells remain much closer to one another. Viewed together with Figure 12, Figure 13 suggests that the deeper-tree log-loss advantage of MinimaxSplit is not merely a numerical accident: it is explained by a clear difference in partition geometry.
Figure 13: Leaf-size statistics for the one-dimensional piecewise classification problem. The left panel shows the mean number of samples per leaf, and the right panel shows the standard deviation of the leaf sizes. MinimaxSplit produces terminal cells with much smaller leaf-size dispersion, which is the classification analogue of avoiding highly unbalanced end-cuts.
D.3 Continuous versus atomic covariates
Finally, Figure 14 compares the same piecewise signal in two sampling regimes. In the left panel, $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$ is marginally atomless, corresponding to the setting of Theorem 12. In the right panel, $X$ is sampled from a uniform grid of $41$ equally spaced points in $[0,1]$, which places us in the purely atomic setting of Theorem 13. The plotted quantity is again the expected log-loss.
The qualitative ordering is the same in both regimes. In each case, the CART baselines improve quickly at the first few levels and then deteriorate as the tree continues to refine, whereas the MinimaxSplit curve becomes nearly flat once it reaches its best value. The advantage of MinimaxSplit is somewhat smaller in the atomic experiment than in the atomless one, but it remains visible throughout the deeper levels. This is consistent with the form of (32), where exponential decay in the atomless case is accompanied by an additional discrete-resolution term. Empirically, the atomic curves flatten earlier, which is natural because once the grid resolution and the minimum-leaf constraint become active, all three procedures have limited room for further refinement.
Overall, Figure 14 shows that the classification analogue of MinimaxSplit is not tied to the marginally atomless setting: its advantage in cross-entropy persists when the covariate distribution is discrete, although the attainable gain is moderated by the finite support.
Figure 14: Expected log-loss for the one-dimensional piecewise classification problem under continuous and atomic covariates. Left: $X\mathrel{\mathop{\kern 0.0pt\sim}\limits^{\mathrm{law}}}\mathrm{Unif}(0,1)$. Right: $X$ sampled from a uniform grid on $[0,1]$. The same qualitative pattern appears in both settings: entropy-CART and Gini-CART improve fastest at very small depths, while MinimaxSplit is more stable and attains the best deep-tree surrogate risk.
Appendix E Further numerics
E.1 Performance metrics